# Plug-and-Play Dramaturge: A Divide-and-Conquer Approach for Iterative Narrative Script Refinement via Collaborative LLM Agents

### v1

@@ Introduction
Baseline Methods
Currently, there are no specialized methods focused on script refinement for comparison. We select six strong mainstream LLM models without employing the divide-and-conquer strategy proposed in this paper as baselines. These baselines include: GPT-4o (OpenAI’s general-purpose model), DeepSeek-R1 (specialized in logical reasoning), Qwen-max (Alibaba’s flagship model), GPT-4.1 (OpenAI’s flagship model), Gemini-2.5-pro (Google’s latest model), and GPT-4.1-mini (GPT-4.1’s efficient version). All experiments are conducted using officially released APIs.
@@
Evaluation
Following established practices in existing work, we employ LLM-based evaluation for script quality assessment (Wu et al. 2024; Han et al. 2024; Ma, Qiao, and Liu 2024; Chen et al. 2024b). We observe that directly evaluating long full scripts faces challenges due to the context understanding limitations of LLMs, making it difficult to effectively capture both global structure and details. Moreover, a script may be structurally sound at the global level but weak in local details, or vice versa. Therefore, we design a complementary evaluation framework that enables a comprehensive assessment of both overall quality and scene-level details.
Script-Level Overall Evaluation
@@
Scene-Level Comparative Evaluation
A pairwise comparison is conducted to assess scene-level details of the scripts. The evaluator performs a comparative analysis between the two scripts by traversing corresponding scenes. The analyses across all scenes are then aggregated to produce a final score. To mitigate position bias(Wang et al. 2024), we implement balanced position calibration by randomly alternating the presentation order of compared scenes. This metric highlights granular improvements that may be overlooked in global assessments.
Figure 3: Enhancement in character development. Dramaturge introduces internal conflict and a subplot for Ron, transforming him from a sidekick into a more well-rounded character through foreshadowing and payoff.
Evaluation Metrics
@@
Figure 4: Enhancement in narrative structure. Dramaturge refines storyline to link external conflicts with internal character arcs, introduces subplots, and externalizes abstract themes into concrete plots, creating a more resonant story.
Figure 5: Enhancement of scene presentation. Dramaturge introduces atmospheric intensification and character-environment integration to create a more immersive and psychologically impactful scene, which mirrors characters’ internal states and provides narrative foreshadowing.
Figure 6: Enhancement of dialogue quality. Enhancement of dialogue quality. The improved dialogue incorporates psychological depth through internal thoughts, grounds character relationships in shared history, and uses subtext to create more natural and emotionally compelling conversations.
Results and Analysis
Quantitative Results
@@
Script-Level Overall Evaluation Results
Dramaturge achieves a total score of 87.70, representing a substantial 53.4% improvement over the original scripts (57.18). Despite using a much weaker backbone GPT-4.1-mini, our method significantly outperforms all baseline models and surpasses the strongest baseline Gemini-2.5-pro by 8.3%, demonstrating a greater advantage from architectural design than merely increasing model capacity. This success stems directly from our divide-and-conquer strategy, where top-down information flow ensures that high-level directives systematically guide local modifications, maintaining global narrative consistency and preventing new contradictions.
Table 2: The results present the mean scores and corresponding standard errors (SE) for both script-level overall evaluation and scene-level comparative evaluation.
@@
Original Scripts 57.18 0.80 56.50 1.65
GPT-4o 60.66 0.67 54.12 0.88
Qwen-max 60.26 0.59 54.28 0.81
Deepseek-r1 68.74 0.79 66.24 0.97
GPT-4.1-mini 71.28 0.69 63.92 1.00
GPT-4.1 73.34 0.53 65.00 1.93
Gemini-2.5-pro 80.98 0.57 78.56 0.97
Ours 87.70 0.34 94.20 0.51
Scene-Level Comparative Evaluation Results
Dramaturge achieves 94.20, representing a 66.7% improvement over the original scripts and outperforming all baseline methods, including the strongest Gemini-2.5-pro by 19.9%. Evaluation results indicate that our framework yields greater improvements in granular refinement at the scene level than at the script level. This advantage stems from the SLR stage, where multi-agent reviews are systematically integrated and routed to appropriate locations, enabling precise identification and correction of detail flaws. Several baseline models even exhibit performance drops due to introducing new inconsistencies during revision, which highlights the inherent difficulty of our task.
Statistical Analysis of the Quantitative Results
We conduct a comprehensive statistical analysis of the quantitative evaluation results, including score distributions and standard errors (SE). As shown in Figure 7, our method achieves the highest median scores with the most concentrated distributions, reflecting superior performance and stability. Table 2 further supports this finding, reporting that our method not only obtains the highest average scores but also the lowest SEs across both evaluation levels. The smaller SEs indicate that the statistical performance estimates of our method across the dataset are more precise and statistically reliable, thus providing stronger evidence of its superiority over baseline approaches.
Figure 7: Distribution of scores for script-level overall evaluation and scene-level comparative evaluation across all datasets. The violin plots show the probability density of scores, with the median (dark red lines), first quartile Q1 (blue lines), and third quartile Q3 (yellow lines) marked.
Qualitative Results
@@
Character Development Enhancement
@@
As shown in Figure 3, Ron Weasley evolves from a simple sidekick through: (1) articulation of internal conflicts, (2) introduction of a subplot, and (3) culmination in heroic action that pays off the foreshadowing, creating a well-rounded character arc. Dramaturge transforms one-dimensional characters into psychologically complex figures.
Narrative Structure Enhancement
@@
As shown in Figure 4, the scene presentation is enhanced by atmospheric intensification and character-environment integration. Atmospheric intensification is achieved through environmental personification and heightened sensory details, creating a palpable sense of tension. Character-environment integration applies physical environmental settings to mirror the characters’ internal states. For instance, psychological externalization is used when shadows become metaphors for Harry’s fears. This emotional mirroring forges a strong link between the physical environmental setting and the characters’ emotional journeys, demonstrating Dramaturge’s capability of transforming functional descriptions into immersive and psychologically resonant scenes.
Figure 8: Ablation study shows the effectiveness of multi-agent collaboration for feature decomposition in each stage.
Scene Presentation Enhancement
As shown in Figure 5, the improved dialogue incorporates greater psychological depth by including characters’ internal thoughts. For example, by adding Voldemort’s reflection on his mother’s love, the dialogues reveal hidden motivations and emotional complexities, enriching his portrayal beyond that of a stock villain. The dialogue also grounds the characters’ relationships by referencing past events, making the characters’ interactions feel more genuine and earned. By infusing dialogue with subtext, psychological nuance, and specific character history, Dramaturge produces conversations that are more natural and emotionally compelling.
Dialogue Quality Enhancement
@@
As shown in Figure 6, Dramaturge elevates dialogue from simple exposition to a multi-faceted tool for character revelation and plot advancement. The improved dialogue incorporates greater psychological depth by including characters’ internal thoughts (e.g., Voldemort’s reflection on his mother’s love), revealing hidden motivations and complexities that enrich his portrayal beyond a stock villain. The dialogue also grounds the characters’ relationships in shared history by referencing past events, adding authenticity and weight to their bond. This technique makes the characters’ interactions feel more genuine and earned. By infusing dialogue with subtext, psychological nuance, and specific character history, Dramaturge produces conversations that are more natural, emotionally resonant, and effective at driving the narrative forward.
The complete original and refined script examples are provided in Appendix E.
Ablation Study
@@
Table 3: Ablation study results show incremental improvement by adding task components. Bold indicates significance with p<0.05p<0.05.
@@
Original Scripts 57.18 56.50
+ HCR 70.76 75.47
+ HCR & SLR 74.83 77.38
+ HCR & GR 77.38 81.05
+ HCR & GR & SLR (Full Model) 87.70 94.10
The Influence of Hierarchical Task Architecture
@@
The Influence of Multi-Agent Collaboration
Figure 8 shows that performance improves progressively with the introduction of specialized agents in each stage. It proves the superiority of multi-agent configurations over single-agent baselines with even partial configurations, which confirms the effectiveness of our feature decomposition.
Figure 9: Our method shows significant and continuous improvement across iterations, surpassing all baseline models.
Iterative Improvement Analysis
As shown in Figure 9, we tracked the score changes during the modification process for the script presented in Appendix E to demonstrate our method’s capability for continuous script quality improvement. Our method starts at 58.0 and reaches 90.0 after six iterations, producing a coarse-to-fine refinement: as the improvement margin decreases, the scores gradually converge. In contrast, baseline methods fail to sustain continuous improvement after only one or two iterations. This occurs because uncoordinated modifications will introduce new inconsistencies that lead to score degradation and iteration stop. Our approach first modifies and stabilizes the global narrative structure before refining local details, preventing new structural conflicts, and enabling continuous improvement.
Conclusion
We introduce Dramaturge, a plug-and-play narrative refinement method through a task and feature oriented divide-and-conquer strategy. The iterative framework employs Global Review, Scene-level Review, and Hierarchical Coordinated Revision to achieve coordinated modifications across multiple granularities and locations while maintaining narrative consistency. Experiments show that Dramaturge achieves substantial improvements over original scripts by 53.4% (script-level) and 66.7% (scene-level), outperforming the strongest baseline by 8.3% and 19.9%, respectively. In future work, we will explore controllable script revision in human-AI interactions. Moreover, developing a more systematic and robust evaluation framework to assess narrative quality at multiple levels would be highly valuable for supporting script generation.
Acknowledgments
This paper was sponsored by the CCF-Tencent Rhino-Bird Open Research Fund, and supported in part by the National Natural Science Foundation of China under Grants 62306319.
Appendix
@@ Acknowledgments
Appendix A Appendix A: LLM prompts for Global Review Stage
@@ 7. DO NOT provide more than 2 suggestions
Table 4: Prompt used for the Engagement Evaluator.
Character Evaluator
@@ 6. DO NOT provide more than 2 suggestions
Table 5: Prompt used for the Character Evaluator.
Theme Evaluator
@@
Table 6: Prompt used for the Theme Evaluator.
Narrative Evaluator
@@
Table 7: Prompt used for the Narrative Evaluator.
Global Review Integrator
The Global Review Integrator bridges the high-level recommendations of the Global Review with the newly generated scene sequence, mapping abstract strategies to specific scenes, which ensures that the improvement strategies aligned with the overall narrative structure.
@@
1. Analyze all inspector suggestions for potential conflicts or contradictions
@@ 2. Identify and eliminate duplicate or overlapping recommendations
3. Resolve any contradictions by prioritizing based on scene context and narrative importance
@@ 7. Maintain consistency with the overall scene’s narrative function
- Resolve timing and pacing conflicts between different inspector recommendations
@@
Table 8: Prompt used for the Global Review Integrator.
Appendix B Appendix B: LLM prompts for Scene-level Review Stage
@@ 6. Emotional impact and tension in dialogue
Table 9: Prompt used for the Dialogue Inspector.
Plot Inspector
@@ 8. Weak narrative points
Table 10: Prompt used for the Plot Inspector.
Character Inspector
@@ 5. Character relationships consistency
Table 11: Prompt used for the Character Inspector.
Scene Description Inspector
@@ 6. Balance between description and action
Table 12: Prompt used for the Scene Description Inspector.
Scene-level Review Integrator
The Scene-level Review Integrator synthesizes and consolidates suggestions from multiple specialized agents (Dialogue Inspector, Plot Inspector, Character Inspector, and Scene Description Inspector) into coherent, prioritized recommendations for each scene.

### v2

@@ 4.1.2. Dataset
4.2. Baseline Methods
Currently, there are no specialized methods focused on script refinement for comparison. Thus, we select six strong flagship LLM models without employing the divide-and-conquer strategy proposed in this paper as baselines. These baselines include: GPT-4o (OpenAI’s general-purpose model), DeepSeek-R1 (specialized in logical reasoning), Qwen-max (Alibaba’s flagship model), GPT-4.1 (OpenAI’s flagship model), Gemini-2.5-pro (Google’s latest model), and GPT-4.1-mini (GPT-4.1’s efficient version). All experiments are conducted using officially released APIs.
@@ 4.2. Baseline Methods
4.3. Evaluation
Following established practices in existing work, we employ LLM-based evaluation for script quality assessment (Wu et al., 2024; Han et al., 2024; Ma et al., 2024; Chen et al., 2024b; Chang et al., 2024; Lee et al., 2025). We observe that directly evaluating long full scripts faces challenges due to the context understanding limitations of LLMs, making it difficult to effectively capture both global structure and details. Moreover, a script may be structurally sound at the global level but weak in local details, or vice versa. Therefore, we design a complementary evaluation framework that enables a comprehensive assessment of both overall quality and scene-level details.
4.3.1. Script-Level Overall Evaluation
@@
4.3.2. Scene-Level Comparative Evaluation
A pairwise comparison is conducted to assess scene-level details of the scripts. The evaluator performs a comparative analysis between the two scripts by traversing corresponding scenes. The analyses across all scenes are then aggregated to produce a final score. To mitigate position bias (Wang et al., 2024), we implement balanced position calibration by randomly alternating the presentation order of compared scenes. This metric highlights granular improvements that may be overlooked in global assessments.
4.3.3. Evaluation Metrics
@@
5. Results and Analysis
5.1. Quantitative Results
@@
5.1.1. Script-Level Overall Evaluation Results
Dramaturge achieves a total score of 84.60, representing a substantial 53.2% improvement over the original scripts (55.24). Despite using a much weaker backbone GPT-4.1-mini, our method significantly outperforms all baseline models and surpasses the strongest baseline Gemini-2.5-pro by 8.2%, demonstrating a greater advantage from architectural design than merely increasing model capacity. This success stems directly from our divide-and-conquer strategy, where top-down information flow ensures that high-level directives systematically guide local modifications, maintaining global narrative consistency and preventing new contradictions.
5.1.2. Scene-Level Comparative Evaluation Results
Dramaturge achieves 90.00, representing a 65.1% improvement over the original scripts and outperforming all baseline methods, including the strongest Gemini-2.5-pro by 18.4%. Evaluation results indicate that our framework yields greater improvements in granular refinement at the scene level than at the script level. This advantage stems from the SLR stage, where multi-agent reviews are systematically integrated and routed to appropriate locations, enabling precise identification and correction of detail flaws. Several baseline models even exhibit performance drops due to introducing new inconsistencies during revision, which highlights the inherent difficulty of our task.
Table 2. The results present the mean scores and corresponding standard errors (SE) for both script-level overall evaluation and scene-level comparative evaluation.
@@ 5.1.2. Scene-Level Comparative Evaluation Results
Original Scripts 55.24 0.78 54.50 1.60
GPT-4o 58.58 0.65 52.20 0.85
Qwen-max 58.18 0.57 52.36 0.78
Deepseek-r1 66.38 0.76 63.96 0.94
GPT-4.1-mini 68.82 0.67 61.70 0.97
GPT-4.1 70.78 0.51 62.72 1.87
Gemini-2.5-pro 78.18 0.55 76.00 0.94
Ours 84.60 0.33 90.00 0.49
5.2. Statistical Analysis of the Quantitative Results
We conduct a comprehensive statistical analysis of the quantitative evaluation results, including score distributions and standard errors (SE). As shown in Fig. 3, our method achieves the highest median scores with the most concentrated distributions, reflecting superior performance and stability. Table 2 further supports this finding, reporting that our method not only obtains the highest average scores but also the lowest SEs across both evaluation levels. The smaller SEs indicate that the statistical performance estimates of our method across the dataset are more precise and statistically reliable, thus providing stronger evidence of its superiority over baseline approaches.
Figure 3. Distribution of scores for script-level overall evaluation and scene-level comparative evaluation across all datasets. The violin plots show the probability density of scores, with the median (dark red lines), first quartile Q1 (blue lines), and third quartile Q3 (orange lines) marked.
5.3. Human Evaluation
We further conducted a blind human evaluation with 42 professional scriptwriters to compare our refined scripts against each baseline across four dimensions: Character Development, Narrative Structure, Dialogue Quality, and Scene Presentation, which are the same dimensions used in the LLM-based evaluation. For every baseline, two disjoint panels of three raters (six raters total) independently judged the assigned script pairs, with each panel responsible for half of the evaluation set so that every pair was assessed by exactly one three-rater panel. Within a panel, raters viewed the pairs and cast individual Win/Tie/Lose votes, and the majority vote produced the script-level judgment. Aggregating these judgments yielded the percentages reported in Table 3.
Table 3. Blind human ratings (Win/Tie/Lose %) comparing Ours against each baseline across four dimensions.
Method Character Development Narrative Structure Dialogue Quality Scene Presentation
Original Scripts 88/8/4 86/10/4 90/7/3 92/6/2
GPT-4o 76/18/6 74/20/6 78/17/5 80/16/4
Qwen-max 74/20/6 73/21/6 76/19/5 78/18/4
Deepseek-r1 68/24/8 66/26/8 70/23/7 72/22/6
GPT-4.1-mini 71/21/8 69/23/8 73/20/7 75/20/5
GPT-4.1 64/26/10 61/28/11 66/25/9 68/24/8
Gemini-2.5-pro 60/30/10 58/31/11 63/28/9 66/26/8
Overall, humans strongly prefer our method across all baselines and dimensions. Against Original Scripts, our method wins 88-92% across dimensions. Against GPT-4o, our approach wins 76-80% across dimensions. Even compared with the strongest baseline Gemini-2.5-pro, our system is preferred in 60-66% of cases with modest tie rates (26-31%) and low loss rates (8-11%). These results corroborate the automatic evaluations, indicating that our divide-and-conquer refinement improves both global structure and local details in ways consistently perceived by human experts.
5.4. Qualitative Results
@@
5.4.1. Character Development Enhancement
@@
As shown in Fig. 4, Ron Weasley evolves from a simple sidekick through: (1) articulation of internal conflicts, (2) introduction of a subplot, and (3) culmination in heroic action that pays off the foreshadowing, creating a well-rounded character arc. Dramaturge transforms one-dimensional characters into psychologically complex figures.
Figure 4. Enhancement in character development. Dramaturge introduces internal conflict and a subplot for Ron, transforming him from a sidekick into a more well-rounded character through foreshadowing and payoff.
5.4.2. Narrative Structure Enhancement
Figure 5 illustrates how Dramaturge enhances narrative architecture through three key mechanisms: converting external obstacles into internal character confrontations, weaving character-specific subplots that generate meaningful interpersonal dynamics, and materializing abstract thematic concepts into tangible story elements. This approach creates narrative cohesion where plot advancement directly serves character evolution.
Figure 5. Enhancement in narrative structure. Dramaturge refines storyline to link external conflicts with internal character arcs, introduces subplots, and externalizes abstract themes into concrete plots, creating a more resonant story.
5.4.3. Scene Presentation Enhancement
@@
As shown in Fig. 6, the scene presentation is enhanced by atmospheric intensification and character-environment integration. Atmospheric intensification is achieved through environmental personification and heightened sensory details, creating a palpable sense of tension. Character-environment integration applies physical environmental settings to mirror the characters’ internal states. For instance, psychological externalization is used when shadows become metaphors for Harry’s fears. This emotional mirroring forges a strong link between the physical environmental setting and the characters’ emotional journeys, demonstrating Dramaturge’s capability of transforming functional descriptions into immersive and psychologically resonant scenes.
Figure 6. Enhancement of scene presentation. Dramaturge introduces atmospheric intensification and character-environment integration to create a more immersive and psychologically impactful scene, which mirrors characters’ internal states and provides narrative foreshadowing.
5.4.4. Dialogue Quality Enhancement
@@
As shown in Fig. 7, Dramaturge elevates dialogue from simple exposition to a multi-faceted tool for character revelation and plot advancement. The improved dialogue incorporates greater psychological depth by including characters’ internal thoughts (e.g., Voldemort’s reflection on his mother’s love), revealing hidden motivations and complexities that enrich his portrayal beyond a stock villain. The dialogue also grounds the characters’ relationships in shared history by referencing past events, adding authenticity and weight to their bond. This technique makes the characters’ interactions feel more genuine and earned. By infusing dialogue with subtext, psychological nuance, and specific character history, Dramaturge produces conversations that are more natural, emotionally resonant, and effective at driving the narrative forward. The complete original and refined script examples are provided in Appendix E.
Figure 7. Enhancement of dialogue quality. The improved dialogue incorporates psychological depth through internal thoughts, grounds character relationships in shared history, and uses subtext to create more natural and emotionally compelling conversations.
5.5. Ablation Study
@@
Table 4. Ablation study results show incremental improvement by adding task components. Bold indicates significance with p<0.05p<0.05.
@@
Original Scripts 55.24 54.50
+ HCR 68.30 72.90
+ HCR & SLR 72.20 74.70
+ HCR & GR 74.70 78.30
+ HCR & GR & SLR (Full Model) 84.60 90.00
5.5.1. The Influence of Hierarchical Task Architecture
@@
5.5.2. The Influence of Multi-Agent Collaboration
Fig. 8 shows that performance improves progressively with the introduction of specialized agents in each stage. It proves the superiority of multi-agent configurations over single-agent baselines with even partial configurations, which confirms the effectiveness of our feature decomposition.
Figure 8. Ablation study shows the effectiveness of multi-agent collaboration for feature decomposition in each stage.
5.6. Iterative Improvement Analysis
[…]
Figure 9. Our method shows significant and continuous improvement across iterations, surpassing all baseline models.
5.7. Computational Efficiency
Our method is designed foremost to improve script quality and, to our knowledge, is the first to deliver stable, iteration-by-iteration improvements for long-form script refinement. Although Dramaturge performs multiple passes of reading and revision, most operations are localized rather than full-script processing, so the incremental overhead per pass is limited. Consequently, the per-iteration LLM invocation cost is only approximately 51% higher than single-pass generation methods such as Critics (Bae and Kim, 2024). In terms of actual runtime, agents within the GR, SLR, and HCR stages execute in parallel, rendering the time of a single iteration comparable to ordinary generation. In practice, the procedure typically converges within about six iterations, keeping the overall cost controllable.
6. Conclusion
We introduce Dramaturge, a plug-and-play narrative refinement method through a task and feature oriented divide-and-conquer strategy. The iterative framework employs Global Review, Scene-level Review, and Hierarchical Coordinated Revision to achieve coordinated modifications across multiple granularities and locations while maintaining narrative consistency. Experiments show that Dramaturge achieves substantial improvements over original scripts by 53.2% (script-level) and 65.1% (scene-level), outperforming the strongest baseline by 8.2% and 18.4%, respectively. In future work, we will explore controllable script revision in human-AI interactions. Moreover, developing a more systematic and robust evaluation framework to assess narrative quality at multiple levels would be highly valuable for supporting script generation.
Appendices
@@ 6. Conclusion
Appendix A LLM prompts for Global Review Stage
@@ 7. DO NOT provide more than 2 suggestions
A.2. Character Evaluator
@@ 6. DO NOT provide more than 2 suggestions
A.3. Theme Evaluator
@@
A.4. Narrative Evaluator
@@
A.5. Global Review Integrator
The Global Review Integrator bridges the high-level recommendations of the Global Review with the newly generated scene sequence, mapping abstract strategies to specific scenes while computing hierarchical priority scores to keep the plan aligned with the overall narrative structure.
Table 9. Prompt used for the Global Review Integrator.
@@ A.5. Global Review Integrator
1. Analyze all evaluator suggestions for potential conflicts or contradictions
@@ 2. Identify and eliminate duplicate or overlapping recommendations
3. Resolve any contradictions by prioritizing based on scope, seriousness, and narrative importance
@@ 7. Maintain consistency with the overall scene’s narrative function
Priority Scoring Protocol:
- Assign each suggestion to one of three scope groups before scoring:
(1) structural directives reshaping arcs/causality across scenes;
(2) cross-scene adjustments affecting character continuity or thematic progression;
(3) localized detail optimizations. Higher-numbered groups never outrank lower-numbered ones.
- Within each scope group, boost suggestions tied to the main plotline and penalize secondary subplots.
- Still within the group, prefer cross-scene or multi-scene effects over single-scene impacts.
- If two suggestions remain tied after the above, break the tie by evaluator order: Narrative, Character, Theme, then Engagement.
@@
- Resolve timing and pacing conflicts between different evaluator recommendations
@@
Appendix B LLM prompts for Scene-level Review Stage
@@ 6. Emotional impact and tension in dialogue
B.2. Plot Inspector
@@ 8. Weak narrative points
B.3. Character Inspector
@@ 5. Character relationships consistency
B.4. Scene Description Inspector
@@
Table 13. Prompt used for the Scene Description Inspector.
@@ 6. Balance between description and action
B.5. Scene-level Review Integrator
The Scene-level Review Integrator synthesizes and consolidates suggestions from multiple specialized agents (Dialogue Inspector, Plot Inspector, Character Inspector, and Scene Description Inspector) into coherent, prioritized recommendations for each scene through a three-stage priority scoring policy aligned with the Global Review directives.
Table 14. Prompt used for the Scene-level Review Integrator.